OpenStudy (anonymous):

QUESTION IS DRIVING ME INSANE!! Help please: Vector U = (4,-4,-3) and Vector V = (5,4,-1) Find: (U dot V) U I found that it's (28,-28,21) but I'm told that's wrong. What did I do wrong??

OpenStudy (anonymous):

@Jack1

OpenStudy (jack1):

g'day man, howzit?

OpenStudy (anonymous):

It's not good Jack1. This question is annoying me more than a scratched tongue.

OpenStudy (jack1):

you're like 99% there dude... negative x negative x negative = negative so (28,-28,-21) not 21...

OpenStudy (jack1):

all better now ;)

OpenStudy (anonymous):

But 7(4,-4,-3) how 21?

OpenStudy (jack1):

gotta go dude, lectures bout to start

OpenStudy (jack1):

will post in class, sry

OpenStudy (anonymous):

Fine, thanks and good luck.

OpenStudy (anonymous):

Oh, for when you get back, I did have -21, it was a typo when I pute "21". So something is still amiss.

OpenStudy (anonymous):

Please help :p

OpenStudy (eamier):

i think it is correct if you get like jack1 said

OpenStudy (anonymous):

That would seem to be wrong though, I do my HW online. So it's both wrong.

OpenStudy (anonymous):

And I did actually get it like him the first time around. Still wrong.

OpenStudy (eamier):

sorry then, looking forward for answer too.

OpenStudy (anonymous):

:(

OpenStudy (jack1):

sorry minato: your dot product (U.V) will give you a vector magnitude only, as you did above so U.V = Ux * Vx + Uy + Vy + Uz + Vz = 7 overall then you are multiplying that by a vector: U, so the answer will be a vector so U.V times U = 7 times (4,-4,-3) = (7 * 4) , (7 * -4) , (7 * -3) = 28, -28, -21 so final answer = (28, -28, -21) = a vector, rather than how the dot product was just a magnitude

OpenStudy (jack1):

just read that it was a typo... I believe the answer u and I got is correct man